### v2

# How to Tell When a Result Will Replicate: Significance and Replication in Distributional Null Hypothesis Tests

Abstract
There is a well-known problem in Null Hypothesis Significance Testing: many statistically significant results fail to replicate in subsequent experiments. We show that this problem arises because standard ‘point-form null’ significance tests consider only within-experiment but ignore between-experiment variation, and so systematically underestimate the degree of random variation in results. We give an extension to standard significance testing that addresses this problem by analysing both within- and between-experiment variation. This ‘distributional null’ approach does not underestimate experimental variability and so is not overconfident in identifying significance; because this approach addresses between-experiment variation, it gives mathematically coherent estimates for the probability of replication of significant results. Using a large-scale replication dataset (the first ‘Many Labs’ project), we show that many experimental results that appear statistically significant in standard tests are in fact consistent with random variation when both within- and between-experiment variation are taken into account in this approach. Further, grouping experiments in this dataset into ‘predictor-target’ pairs we show that the predicted replication probabilities for target experiments produced in this approach (given predictor experiment results and the sample sizes of the two experiments) are strongly correlated with observed replication rates. Distributional null hypothesis testing thus gives researchers a statistical tool for identifying statistically significant and reliably replicable results. Keywords: Hypothesis Testing; Replication; Significance;
Confronted with a surprising experimental result and wondering whether it may represent a meaningful discovery, a researcher must ask: could this result simply be a consequence of random variation and no real effect? Such questions are especially important for discovery in domains of high variability (psychology, neuroscience, medicine, genetics, and so on), where random variation can easily produce apparently surprising results. Since by definition a surprising result has not been hypothesised a priori, such questions cannot be answered by tests which compare hypotheses with distinct a priori means (e.g. Neyman-Pearson hypothesis testing or informed Bayesian model comparison; Christensen,, 2005; Berger and Pericchi,, 1996; Gönen et al.,, 2005). The standard way to answer such questions is via Fisher’s evidential significance test, which assesses the degree to which a result is unexpected under the null statistical model of random variation but no real effect (and so gives evidence against that model) operationalised as the probability $p$ of a result as or more extreme arising under that model.
While the logic of evidential significance testing is compelling (a result is surprising when it is extreme, relative to our expectations; a low $p$-value tells us that our result is extreme, relative to the null model) it has become increasingly clear that significance testing, at least as it is typically used, is fundamentally flawed. Perhaps the most important problem with standard significance testing involves replication and the reliability of results, with standard significance testing being overconfident in identifying apparently nonsensical effects, such as telepathy or precognition, as real (parapsychology ‘offers a truly alarming massive case study of how statistics can mislead and be misused’: Diaconis et al.,, 1991), and more generally, with systematic research showing that many statistically significant experimental results in psychology and related fields do not occur reliably in replications (e.g. Camerer et al.,, 2018; Open Science Collaboration et al.,, 2015; Klein et al.,, 2018, 2014), results seen as ‘reflecting an unprecedented level of doubt among practitioners about the reliability of research findings in the field’ (Pashler and Wagenmakers,, 2012). These flaws are also evident in well-known effects of sample size on significance (the observation that the probability of getting a statistically significant result in the standard approach increases with sample size, irrespective of the presence or absence of a true effect; to quote Thompson, (1998): ‘Statistical testing becomes a tautological search for enough participants to achieve statistical significance. If we fail to reject, it is only because we’ve been too lazy to drag in enough participants’) and in the fact that in standard significance tests, null hypotheses are always false (and to quote Cohen, 2016: ‘if the null hypothesis is always false, what’s the big deal about rejecting it?’). Faced with these problems and others (primarily related to hypothesis comparison and optional stopping, e.g. Wagenmakers et al.,, 2018; Wagenmakers,, 2007) there have been increasingly widespread calls for the wholesale abandonment of the significance testing approach (e.g. McShane et al.,, 2019; Amrhein and Greenland,, 2018; Wagenmakers et al.,, 2018; Hunter,, 1997; Carver,, 1978).
Our aim in this paper is to show that these problems all arise from a basic gap in the standard approach to significance testing: the fact that it considers only within-experiment random variation in estimation, while ignoring between-experiment random variation. In the language of measurement analysis in physics, the standard approach considers the precision of the experiment as a measurement tool (the degree of random variation in measurements within this experiment, a factor which depends on sample size) but ignores the accuracy of the experiment as a tool (the degree to which measurements in this experiment agree with those obtained from other experiments measuring the same effect; e.g Mandel,, 2012). This gap means that the standard approach systematically underestimates the degree of random variation or error affecting experimental results, and so is overconfident in identifying results as significant. We address this by describing a model of significance testing that includes both within- and between-experiment random variation; the chance of statistical significance in this ‘distributional’ model does not rise with sample size irrespective of any real effect; this approach does not overestimate the random variation affecting experimental results and so is not overconfident in identifying significance; importantly, this model gives coherent estimates for the probability of replication of significant results in future experiments.
The model we describe mixes statistical inference typically associated with the Frequentist approach with that typically associated with the Bayesian: in section $1$ we give a general sketch and overview of the model, positioning it relative to these two approaches. In section $2$ we derive basic expressions for significance, statistical power and probability of replication in this model in the context of a $t$-test, assuming known between-experiment variation. In section $3$ we give unbiased sample estimators for between-experiment variance of means, give expressions for significance and replication in terms of those sample estimates, and extend these to give simple and practical expressions for significance and replication in terms of the standard $T$ distribution. We also extend this approach to tests of correlation and linear regression, and tests of contingency, and briefly compare this model to a well-known point-form approach to replication (Killeen, 2005a, ; Killeen,, 2007). In section $4$ we test this model against the first ‘Many Labs’ replication dataset (Klein et al.,, 2014), showing that this model’s assumption of between-experiment variation holds, that its judgements of statistical significance are many orders of magnitude more conservative than those obtained in the standard approach, and that its replication probabilities reliably and accurately predict observed rates of successful replication of individual results. In section $5$ we compare this model to Bayesian approaches to hypothesis testing via the JZS Bayesian $t$-test (e.g. Rouder et al.,, 2009), showing that this Bayesian approach gives results almost identical to those of a point-form significance test, at least in the Many Labs dataset. In the final section we address some possible criticisms of this model of significance and replication, and argue that this distributional approach to evidential significance testing will help researchers identify experimental results which are both surprising (statistically significant) and amenable to further investigation (replicable). Rather than abandoning significance testing as a way of identifying surprising and interesting results, researchers should consider using significance testing in its more general distributional form.
1 Overview of the distributional model
The basic distinction between Frequentist and Bayesian approaches to statistical inference lies in the interpretation of probability: the Frequentist sees probability as the limit of relative frequency in sampling, while the Bayesian sees probability as representing degree of belief or uncertainty in estimates. This distinction means that Frequentist statistical inference considers the distribution of measurements in samples, while Bayesian inference considers both this sampling distribution and additional prior distributions (with these priors reflecting degrees of belief or uncertainty in parameters of the sampling distribution). For example: in a simple experiment estimating some unknown parameter $\mu$ from $N$ measurements and assuming Normal error with variance $\sigma^{2}$, the standard Frequentist approach assumes that the sample mean $\overline{X}$ is distributed as
$\overline{X}\sim\mathcal{N}(\mu,\sigma^{2}/N)$
and, assuming for simplicity that the sample variance $S^{2}$ is an accurate estimate of $\sigma^{2}$, a given result is statistically significant relative to the null hypothesis $H_{0}:\mu=0$ at level $\alpha$ when
$P(\overline{X}|H_{0})=2\Phi\left(-\frac{|\overline{X}|}{S/\sqrt{N}}\right)<\alpha$
where $P(\overline{X}|H_{0})$ represents the probability of observing a result equal to or more extreme than $\overline{X}$ given $H_{0}$, and $\Phi$ represents the cumulative Normal distribution.
A Bayesian approach in this situation, by contrast, assumes a prior distribution of $\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$ with $\mu_{0}$ representing an estimate for the parameter of interest and $\sigma_{0}^{2}$ representing uncertainty or degree of belief in that estimate, so that
$\overline{X}\sim\mathcal{N}(\mu,S^{2}/N)\textit{\, where \,}\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})\textit{\, }$
or simplifying by integrating out $\mu$
$\overline{X}\sim\mathcal{N}(\mu_{0},\sigma^{2}/N+\sigma_{0}^{2})$
and so taking $H_{0}:\mu_{0}=0$ (and assuming again that $S^{2}$ accurately estimates $\sigma^{2}$) this gives
$P(\overline{X}|H_{0})=2\Phi\left(-\frac{|\overline{X}|}{\sqrt{S^{2}/N+\sigma_{0}^{2}}}\right)$
In fact, however, Bayesians are typically more interested in the updated or posterior distribution
$\mu\sim\mathcal{N}(\mu_{N},\sigma_{N}^{2})$
which is obtained by applying Bayes’ theorem to the prior and the $N$ sample measurements. This gives a probability distribution for hypotheses $H$ given data $\overline{X}$ of
$P(H|\overline{X})=P(\mu=m|\overline{X})=\mathcal{N}(m|\mu_{N},\sigma_{N}^{2})$
which, in the Bayesian perspective, describes how we should update our beliefs about $\mu$ given the observed data $\overline{X}$.
The standard Frequentist approach considers only within-experiment variance of the mean, $\sigma^{2}/N$, but ignores any between-experiment variance in means. Our distributional model extends the Frequentist approach to include between-experiment variance of means, which we write as $\sigma_{0}^{2}$. This gives a distribution for means of
$\overline{X}\sim\mathcal{N}(\mu_{0},S^{2}/N+\sigma_{0}^{2})$
just as in the Bayesian approach. Given this distribution, and estimating $\sigma_{0}^{2}$ by the sample variance or heterogeneity of means across similar experiments, $S^{2}_{0}$ (just as $\sigma^{2}$ is estimated by the sample variance of measurements within a single experiment, $S^{2}$) we can assess statistical significance relative to both within- and between-experiment variance: a given result will be statistically significant relative to the null hypothesis $H_{0}:\mu_{0}=0$ when
$P(\overline{X}|H_{0})=2\Phi\left(\frac{-|\overline{X}|}{\sqrt{S^{2}/N+S^{2}_{0}}}\right)<\alpha$
We can also produce a posterior distribution $\mu\sim\mathcal{N}(m|\mu_{N},\sigma_{N}^{2})$ by applying Bayes’ theorem as before, giving a distribution for hypotheses given data of $P(H|\overline{X})$. In this distributional approach, both directions of inference (probability of data given hypothesis $P(\overline{X}|H)$ and probability of hypotheses given data $P(H|\overline{X})$) can be applied within the Frequentist perspective, with the difference between Frequentist and Bayesian viewpoints arising in the interpretation of $\sigma_{0}^{2}$ (the Frequentist interpreting it as representing between-experiment variance of experimental means and the Bayesian as representing degree of prior belief or uncertainty about the population mean). Given the same experimental result $\overline{X}$ and the same values for $\mu_{0}$ and $\sigma_{0}^{2}$, however interpreted, in this approach Frequentists and Bayesians will make the same judgement of statistical significance and will infer the same updated distribution for $\mu$.
Our position, then, is that Bayesian and Frequentist approaches to statistical inference differ in their interpretation of probability, but not in the analyses that can be applied to mathematical representations of probability in terms of statistical distributions. It happens that analyses which estimate $P(\overline{X}|H)$ are primarily associated with the Frequentist viewpoint (most likely because estimates $P(H|\overline{X})$ are uninformative under a point-form null hypothesis: when the probability density of $H$ is massed at a single point, $P(H|\overline{X})=P(H)$ necessarily holds) while those which estimate $P(H|\overline{X})$ are primarily associated with the Bayesian approach.111 Although some Bayesian proposals do estimate $P(\overline{X}|H)$, typically as part of a prior check of validity before Bayesian inference (e.g. Dass and Berger,, 2003; Gabry et al.,, 2019; van de Schoot et al.,, 2021; Schad et al.,, 2021). Our distributional model, by contrast, makes use of both $P(\overline{X}|H)$ and $P(H|\overline{X})$ throughout analysis, to estimate both the significance of new results and the probability of obtaining similar results in future replications.
Some readers may initially see this model as a confusing or contradictory mix of Bayesian and Frequentist interpretations of probability, a perception arising from the fact that the model involves both inference from hypothesis to data (typically Frequentist) and inference from data to hypothesis (typically Bayesian). This perception is incorrect, because these two forms of conditional inference apply to mathematical representations of probability as statistical distributions, not to interpretations of probability (as limiting relative frequencies or as degrees of belief). As such these two forms of inference can be applied simultaneously, with no contradiction, under any interpretation where probabilities are represented as distributions. We expand this initial sketch to give a full and detailed presentation of our distributional model below.
2 Significance and replication for the $t$ distribution
We consider an experiment involving $N$ measurements which we assume follow a Normal distribution
$X\sim\mathcal{N}(\mu,\sigma^{2})$
with unknown mean $\mu$ and unknown variance $\sigma^{2}$ (the standard assumption of normally distributed error), so that the sample mean $\overline{X}$ follows the distribution $\mathcal{N}(\mu,\sigma^{2}/N)$. Given sample variance $S^{2}$ and degrees of freedom $\nu$, the variance in $\overline{X}$ is estimated by $S^{2}/N$ (the sample variance of the mean) and the standard approach assesses significance against the point-form null hypothesis $\mu=0$ via a $t$-test.
Letting $T_{\nu}$ represent the cumulative $t$ distribution with degrees of freedom $\nu$, $d=\overline{X}/S$ the normalised sample mean and $t=d\sqrt{N}$ the test statistic, in this approach the point-form significance of a result $t$ in a two-sided test is $p=2T_{\nu}(-|t|)$. Since $d$ estimates the true normalised mean $\mu/\sigma$ (a constant), the expected value of $|t|$ rises with $N$ when $\mu\neq 0$, and so the expected value of $p$ falls with $N$ to a limit of 00. Since $\mu=0$ can never hold exactly, in the standard approach to significance testing every experiment is expected to achieve significance for large enough $N$.
The sample variance of the mean in a given experiment is a function of the degree of random variation within that experiment. This within-experiment variation arises as a consequence of random factors affecting individual measurements in that experiment, with each factor moving individual responses randomly one way or another. If the number of such random confounding factors is large we can, via the central limit theorem, assume that the combined effect of all these factors will follow a Normal distribution (irrespective of the statistical distribution of the individual factors themselves), so that the overall within-experiment variation is approximately Normal.
This within-experiment variation is, however, not the only form of variation in experimental means: means are also affected by between-experiment variation (random variation in experimental means from one experiment to another). This between-experiment variation arises as a consequence of factors that have a fixed or systematic effect within a given experiment, but vary randomly across different experiments with each factor moving the mean $\mu$ in a given experiment randomly one way or another. Assuming that the number of such random between-experiment factors is also large, the overall between-experiment variation is, again, approximately Normal. Between-experiment variation is very familiar to researchers: a core aim in experimental design is to control for factors or confounds that may influence results. Such factors, however, can never be fully controlled (no experiment is perfect) and so necessarily cause some degree of random variation in experimental results: variation which should be taken into account when assessing significance.
Since the standard approach to significance testing only considers within-experiment variation (sample variance of the mean) it systematically underestimates the degree of random variation in results, and so will identify results as ‘statistically significant’ when in fact they arise as a consequence of random between-experiment variation. Our model assumes that $\mu$ itself varies randomly across experiments following the distribution
$\mu\sim\mathcal{N}(\mu_{0},\sigma_{0}^{2})$
with unknown variance $\sigma_{0}^{2}$, with the null hypothesis being $\mu_{0}=0$. The variation of $\overline{X}$ around the overall mean $\mu_{0}$ now has two components: within-experiment variance $\sigma^{2}/N$ and between-experiment variance $\sigma_{0}^{2}$. As $N$ increases the within-experiment variance of the mean, $\sigma^{2}/N$, falls and between-experiment variance comes to be the dominant cause of random variation in means.
2.1 Significance for a given variance ratio
We characterise the degree of between-experiment variance $\sigma_{0}^{2}$ via the variance ratio $b=\sigma_{0}^{2}/\sigma^{2}$. For a given value of $b$ we have
$p(\overline{X}|N,\sigma,b)=\int_{-\infty}^{\infty}\mathcal{N}(\overline{X}|\mu,\,\sigma^{2}/N)\mathcal{N}\left(\mu|\mu_{0}=0,b\sigma^{2}\right)\mathrm{d}\mu=\mathcal{N}\left(\overline{X}|0,\,(\sigma^{2}/N)(1+bN)\right)$
and so
$\frac{\overline{X}}{(\sigma/\sqrt{N})\sqrt{1+bN}}$
is normally distributed with mean 00 and variance $1$. Letting
$t=\frac{\overline{X}}{\sqrt{S^{2}}}\sqrt{N}$
as before, we thus see that
$\frac{\frac{\overline{X}}{(\sigma/\sqrt{N})\sqrt{1+bN}}}{\sqrt{(\frac{\nu\ S^{2}}{\sigma^{2}})/\nu}}=\frac{t}{\sqrt{1+bN}}$
follows a $t$-distribution with $\nu$ degrees of freedom, and so the statistical significance of a result $t$ in a two-sided test relative to the variance ratio $b$ is
$p_{sig}(t|b)=2T_{\nu}\left(-|t|/\sqrt{1+bN}\right)$ (1)
Letting $t_{crit}=T^{-1}_{\nu}(1-\alpha/2)$ be the critical value for significance in a two-sided test, a result $t$ will thus be significant at level $\alpha$ when
$t_{crit}<\frac{|t|}{\sqrt{1+bN}}$
or, substituting effect size $d=t/\sqrt{N}$, when
$t_{crit}<\frac{|d|}{\sqrt{b+1/N}}$
Since the $T$ distribution approaches the Normal distribution with rising $N$ and the sample effect size $|d|$ similarly approaches the true effect size $|\mu/\sigma|$ for our experiment, we see that sample results from an experiment where
$\frac{|\mu/\sigma|}{\sqrt{b}}<\Phi^{-1}\left(1-\alpha/2\right)\,\,\,\rightarrow\,\,\,|\mu/\sigma_{0}|<\Phi^{-1}\left(1-\alpha/2\right)$
are never expected to reach significance at level $\alpha$, irrespective of sample size $N$ and, unlike with a point-form null, a distributional null hypothesis is not ‘always false’.
2.2 Significance, $\alpha$ and type $1$ error rate
The $p$ (in a point-form analysis) or $p_{sig}$ (in a distributional analysis) value of a given experimental result tells us the probability of a result as or more extreme arising under our (point-form or distributional) null hypothesis. Given this probability a researcher must decide whether to treat this result as significant. In Fisher’s test for significance this decision is made by comparing the obtained probability against some subjective criterion $\alpha$, such that a probability less than $\alpha$ is counted as significant. This criterion $\alpha$ represents the type 1 error rate of the test; that is, the probability of incorrectly rejecting the null hypothesis when that hypothesis is true.
Note that that this link between $\alpha$ and error rate holds independently of the form of the hypothesis: a point-form test $p<\alpha$ will incorrectly reject the point-form null hypothesis at rate $\alpha$, while a distributional test $p_{sig}<\alpha$ will incorrectly reject the distributional null hypothesis at rate $\alpha$. This does not, however, mean that point-form and distributional tests at some level $\alpha$ are in any way equivalent. On the contrary, the fact that the distributional approach takes between-experiment variation into account while the point-form approach does not means that many results for which point-form $p<\alpha$ (suggesting a significant result) will simultaneously have $p_{sig}>\alpha$ (the result is not significant). Indeed, the difference between point-form and distributional significance can be extremely large: in our analysis of experiments in the Many Labs $1$ dataset (below), we found that around $10\%$ of experimental results with point-form $p<0.00001$ (highly significant relative to the point-form null) had $p_{sig}>0.05$ (not significant relative to the distributional null).
2.3 Significance and effect size
The size of an effect in a $t$-test is typically estimated in units of sample standard deviation, giving an effect-size estimate of $d=\overline{X}/S$ (an estimate of the true effect $\delta=\mu/\sigma$). Since in the point-form approach every experiment is expected to achieve significance for large enough $N$, effect size is necessarily independent of significance in this approach: the point-form significance of a result tells us nothing about its effect size. Effect size is thus considered separately from significance, with researchers proposing various criteria to distinguish between strong, weak, and negligible effects (Cohen,, 1992; Fritz et al.,, 2012; Kelley and Preacher,, 2012; Cohen,, 2013) alongside various lower bounds on acceptable effect sizes in terms of, for example, experimental ‘crud’ (Meehl, 1990b, ; Meehl, 1990a, ; Orben and Lakens,, 2020).
In the distributional approach, by contrast, significance and effect size are no longer independent. Instead we have
$p_{sig}(t|b)=2T_{\nu}\left(-|t|/\sqrt{1+bN}\right)=2T_{\nu}\left(-|d|/\sqrt{b+1/N}\right)\approx 2T_{\nu}\left(-|d|/\sqrt{b}\right)$
and significance and effect size are directly related: if $p_{sig}(t|b)$ is low then the observed effect $d$ is large relative to the variance ratio $b$, while if $p_{sig}$ is high then the effect $d$ is small. A common recommendation in effect size analyses is to select effect size criteria based on distributions of effect sizes for comparable outcome measures (Lipsey et al.,, 2012). This is just what distributional significance provides: given an estimate of the between-experiment variance ratio $b$ (equivalent to a distribution of effect sizes), in the distributional approach a significance level $\alpha$ places a bound on detectable effects.
2.4 Power, $\beta$ and type $2$ error rate
Where $\alpha$ by construction estimates the type 1 error rate of an hypothesis test (the probability of incorrectly rejecting the null hypothesis when that hypothesis is true), the related measure $\beta$ estimates the type 2 error rate of that test (the probability of failing to reject the null hypothesis when it is in fact false), with $1-\beta$ being the statistical power of the experiment (the probability of correctly rejecting the null hypothesis when it is false). Estimates of $\beta$ are made assuming that some true effect $\delta=\mu/\sigma>0$ holds; here we compare expressions for $\beta$ in point-form and distributional null approaches.
To estimate $\beta$ for tests relative to a point-form null hypothesis, we let $T_{\nu}(x;\theta)$ represent the cumulative probability from $\infty$- ∞ to $x$ of the non-central $T$ distribution with degrees of freedom $\nu$ and non-centrality parameter $\theta$, and assume that $X\sim\mathcal{N}\left(\delta\sigma,\sigma^{2}\right)$ for some effect size $\delta\neq 0$ (so the null is false). In this situation the variable $t=(\overline{X}/S)\sqrt{N}$ follows a non-central $t$-distribution with non-centrality parameter $\delta\sqrt{N}$. The probability, in the point-form approach, of getting a non-significant result in a two-sided test with critical value $t_{crit}$, given that effect $\delta$, is then
$\beta=T_{\nu}\left(t_{crit};|\delta|\sqrt{N}\right)-T_{\nu}\left(-t_{crit};|\delta|\sqrt{N}\right)$
and for an experiment to detect an effect of size $\delta$ with significance $\alpha$ and power $1-\beta$ we must have a sample size $N$ such that
$T_{\nu}\left(t_{crit};|\delta|\sqrt{N}\right)-T_{\nu}\left(-t_{crit};|\delta|\sqrt{N}\right)\leq\beta$
Since
$\lim_{N\rightarrow\infty}T_{\nu}\left(t;|\delta|\sqrt{N}\right)=0$
for all $t$ (as the non-centrality parameter approaches infinity, the area under the non-central $t$ distribution less than any fixed value $t$ necessarily approaches 00) and since
$\begin{split}\lim_{N\rightarrow 0}\left[T_{\nu}\left(t_{crit};|\delta|\sqrt{N}\right)-T_{\nu}\left(-t_{crit};|\delta|\sqrt{N}\right)\right]=T_{\nu}\left(t_{crit}\right)-T_{\nu}\left(-t_{crit}\right)=1-\alpha\end{split}$
(the non-central distribution approaching the central distribution in this situation) we see that in the point-form approach $\beta$ has an upper limit of $1-\alpha$ and falls with increasing $N$ towards a lower limit of 00. The power of a given experiment ($1-\beta$) thus has a minimum of $\alpha$ and rises to a limit of $1$ with increasing $N$ (for any effect $\delta>0$ there is a corresponding sample size $N$ which can detect that effect with any required power).
In the distributional situation we can do a similar analysis in terms of the overall null hypothesis by assuming some true effect $\delta=\mu_{0}/\sigma_{0}>0$, giving
$\overline{X}\sim\mathcal{N}\left(\delta\sigma_{0},(\sigma^{2}/N)(1+bN)\right)$
and so
$\frac{\overline{X}}{(\sigma/\sqrt{N})\sqrt{1+bN}}$
is normally distributed with mean $\delta\sigma_{0}$ and variance $1$, and defining $t$ as before we see that
$\frac{t}{\sqrt{1+bN}}=\frac{\frac{\overline{X}-\delta\sigma_{0}}{(\sigma/\sqrt{N})\sqrt{1+bN}}+\frac{\delta\sigma_{0}}{(\sigma/\sqrt{N})\sqrt{1+bN}}}{\sqrt{\frac{\nu\ S^{2}}{\sigma^{2}}/\nu}}$
follows a non-central $t$-distribution with $\nu$ degrees of freedom and non-centrality parameter
$\frac{\delta\sigma_{0}}{(\sigma/\sqrt{N})\sqrt{1+bN}}=\frac{\delta}{\sqrt{1+1/bN}}$
For a two-sided test we thus have a type $2$ error rate of
$\beta=T_{\nu}\left(t_{crit};\frac{|\delta|}{\sqrt{1+1/bN}}\right)-T_{\nu}\left(-t_{crit};\frac{|\delta|}{\sqrt{1+1/bN}}\right)$
Here the non-centrality parameter falls towards 00 with declining $bN$ and so the error rate $\beta$ rises, approaching an upper limit of $1-\alpha$ as before. As $bN$ rises, $\beta$ falls to a lower limit of
$\begin{split}&\lim_{bN\rightarrow\infty}\left[T_{\nu}\left(t_{crit};\frac{|\delta|}{\sqrt{1+1/bN}}\right)-T_{\nu}\left(-t_{crit};\frac{|\delta|}{\sqrt{1+1/bN}}\right)\right]=T_{\nu}\left(t_{crit};|\delta|\right)-T_{\nu}\left(-t_{crit};|\delta|\right)\end{split}$
We thus see that the power of an experiment rises with $bN$ from a minimum of $\alpha$ towards a maximum of
$1-T_{\nu}\left(t_{crit};|\delta|\right)+T_{\nu}\left(-t_{crit};|\delta|\right)\approx 1-T_{\nu}\left(t_{crit};|\delta|\right)$
(since the last term is negligible) and no experiment can detect an effect $\delta$ with a power greater than this value, irrespective of its sample size $N$ or the variance ratio $b$. Since $T_{\nu}\left(t_{crit};|\delta|\right)>0.5$ when $|\delta|<t_{crit}$ this means that statistical power above $0.5$ can only be achieved for effects $|\delta|>t_{crit}$, and that power approaches $1$ only for very large effect size $|\delta|$. Where in the point-form approach we can design an experiment to detect any effect size at any required level of statistical power (by selecting a large enough $N$) in the distributional approach this is not possible, because the power of an experiment to detect an effect is, in the limit, a function of between-experiment variance rather than sample size.
2.5 Estimating the probability of replication
Because this distributional approach explicitly represents the variation in results across experiments (via the distributional null hypothesis), it allows us to express the probability of obtaining a statistically significant result at level $\alpha$ in a replication of our original experiment. In this section we give estimates for the replication probability, $p_{rep}$, for a given result based first on a fixed value $b$, and then based on a sample estimate $\hat{b}$. In subsequent sections we give simple ‘closed form’ approximations to $p_{sig}$ and $p_{rep}$ in terms of the $T$ distribution (for significance) and the Normal and inverse $\chi^{2}$ distributions (for replication). We also give generic bounds on $p_{sig}$ and $p_{rep}$ that can be used when no estimate for the variance ratio $\hat{b}$ is available.
For some value of $b$ we assume a replication experiment measuring the same $X$ as the original, with sample mean $\overline{X}_{r}\sim\mathcal{N}(\mu_{r},\,\sigma_{r}^{2}/N_{r})$, where $\mu_{r}$ and $\sigma^{2}_{r}$ are the unknown mean and unknown variance of $X$ in the replication, and $N_{r}$ and $\nu_{r}$ the number of measurements and the degrees of freedom. We take $S_{r}^{2}$ to represent the sample variance of $X$ in the replication, and relate within-experiment variance in the original and replication via the ratio $c=\sigma_{r}^{2}/\sigma^{2}$.
Letting $t_{crit_{r}}=T_{\nu_{r}}^{-1}(1-\alpha/2)$ be the critical $t$ value at level $\alpha$ in this replication experiment and
$t_{r}=\frac{\overline{X}_{r}}{\sqrt{S_{r}^{2}}}\sqrt{N_{r}}$
be the test statistic, a result $t_{r}$ will count as a replication of our original result $t$ if it is significant at the same level $\alpha$, and in the same direction, as $t$. In other words, if our original result $t$ was significant in a left-tailed test ($t<-t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}<-t_{crit_{r}}\sqrt{1+bN_{r}}$, while if our original result was significant in a right-tailed test ($t>t_{crit}\sqrt{1+bN}$) then it is replicated if $t_{r}>t_{crit_{r}}\sqrt{1+bN_{r}}$.
To estimate the probability of replication we must derive a distribution for $t_{r}$. We use Bayes’ theorem to update our initial null hypothesis distribution $\mu\sim\mathcal{N}(0,\,\sigma_{0}^{2})$ using the observed value of the sample mean in the original experiment, $\overline{X}$, getting the posterior distribution for $\mu$ of
$\mu\sim\mathcal{N}(\mu_{N},\,\sigma_{N}^{2})$
where
$\sigma_{N}^{2}=\frac{\sigma_{0}^{2}\,\sigma^{2}}{N\,\sigma_{0}^{2}+\sigma^{2}}=\frac{bN}{(1+bN)}(\sigma^{2}/N)$
$\mu_{N}=\frac{\sigma_{N}^{2}}{(\sigma^{2}/N)}\,\overline{X}=\frac{bN}{(1+bN)}\overline{X}$
(a property of the Normal distribution; see e.g. Bishop,, 2006; Murphy,, 2007). Note that this distribution does not represent an unconditional inference about the probability distribution for $\mu$: our experimental result does not tell us that $\mu$ in fact has this distribution (because a distribution for $\mu$ cannot be inferred from a single mean). Instead this is the inferred distribution for $\mu$ conditional on the initial null hypothesis; that is, conditional on the initial assumption that $\mu\sim\mathcal{N}(0,\,\sigma_{0}^{2})$.
Given this updated distribution for $\mu$, the probability density of the sample mean $\overline{X}_{r}$ in our second experiment is
$p(\overline{X}_{r}|\overline{X},b)=\int\mathcal{N}(\overline{X}_{r}|\mu,\,c\sigma^{2}/N_{r})\mathcal{N}(\mu|\mu_{N},\,\sigma_{N}^{2})\mathrm{d}\mu$
and so
$\overline{X}_{r}\sim\mathcal{N}(\mu_{N},\,c\sigma^{2}/N_{r}+\sigma_{N}^{2})\sim\mathcal{N}\left(\frac{bN}{(1+bN)}\,\overline{X},\,(\sigma^{2}/N)\left(\frac{bN}{1+bN}+\frac{cN}{N_{r}}\right)\right)$
and the variable
$\frac{\overline{X}_{r}-\frac{bN}{1+bN}\,\overline{X}}{\frac{\sigma}{\sqrt{N}}\sqrt{\frac{bN}{1+bN}+\frac{cN}{N_{r}}}}$
is normally distributed with mean 00 and variance $1$. This means that the variable
$\frac{\overline{X}_{r}-\frac{bN}{1+bN}\,\overline{X}}{\sqrt{(\frac{\nu_{r}\ S_{r}^{2}}{c\sigma^{2}})/\nu_{r}}\left(\frac{\sigma}{\sqrt{N}}\sqrt{\frac{bN}{1+bN}+\frac{cN}{N_{r}}}\right)}=\frac{t_{r}\sqrt{\frac{cN}{N_{r}}}-\frac{bN}{1+bN}\frac{\overline{X}}{\sqrt{S^{2}_{r}/N}}\sqrt{c}}{\sqrt{\frac{bN}{1+bN}+\frac{cN}{N_{r}}}}$
has a $t$-distribution with $\nu_{r}$ degrees of freedom.
Since the variance in the replication experiment is $c$ times that of the original experiment, the variable $S^{2}_{r}/c$ follows approximately the same distribution as the variable $S^{2}$ (a scaled $\chi^{2}$ distribution) when $N$ and $N_{r}$ are similar (with both variables following exactly the same distribution when $N=N_{r}$), or when both are large. In this situation the variable
$\frac{t_{r}\sqrt{\frac{cN}{N_{r}}}-\frac{bN}{1+bN}\frac{\overline{x}_{1}}{\sqrt{S/N}}}{\sqrt{\frac{bN}{1+bN}+\frac{cN}{N_{r}}}}=\frac{t_{r}\sqrt{\frac{cN}{N_{r}}}-\frac{bN}{1+bN}t}{\sqrt{\frac{bN}{1+bN}+\frac{cN}{N_{r}}}}=\frac{t_{r}\sqrt{c}-t\,\frac{b\sqrt{N\,N_{r}}}{1+bN}}{\sqrt{c+bN_{r}/(1+bN)}}$
will follow a $t$-distribution with $\nu_{r}$ degrees of freedom.
Given this distribution the probability of obtaining a result less than $-t_{crit_{r}}\sqrt{1+bN_{r}}$ in our replication experiment is
$T_{\nu_{r}}\left(\frac{-t_{crit_{r}}\sqrt{c(1+bN_{r})}-t\,\frac{b\sqrt{N\,N_{r}}}{1+bN}}{\sqrt{c+bN_{r}/(1+bN)}}\right)$ (2)
and the probability of obtaining a result greater than $t_{crit_{r}}\sqrt{1+bN_{r}}$ is
$1-T_{\nu_{r}}\left(\frac{t_{crit_{r}}\sqrt{c(1+bN_{r})}-t\,\frac{b\sqrt{N\,N_{r}}}{1+bN}}{\sqrt{c+bN_{r}/(1+bN)}}\right)=T_{\nu_{r}}\left(\frac{t\,\frac{b\sqrt{N\,N_{r}}}{1+bN}-t_{crit_{r}}\sqrt{c(1+bN_{r})}}{\sqrt{c+bN_{r}/(1+bN)}}\right)$ (3)
If our original result $t$ was negative, then the probability of replication is given by Equation 2 while if positive it is given by Equation 3 and so substituting for $t_{crit_{r}}$ we see that the probability of replication of any result $t$ (given variance ratios $b$ and $c$) is
$T_{\nu_{r}}\left(\frac{|t|\frac{b\sqrt{N\,N_{r}}}{1+bN}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)\sqrt{c(1+bN_{r})}}{\sqrt{c+bN_{r}/(1+bN)}}\right)$
Since our second experiment is a replication of the first, we can generally assume that the within-experiment variance $\sigma^{2}$ in both experiments will be approximately the same (so $c=1$) and, simplifying, we have an expression for the probability of replication of
$p_{rep}(t|b)=T_{\nu_{r}}\left(\frac{|t|b\sqrt{N\,N_{r}}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)(1+bN)\sqrt{1+bN_{r}}}{\sqrt{(1+bN+bN_{r})(1+bN)}}\right)$ (4)
2.6 Bounds on significance and replication
These expressions impose useful bounds on significance and replication probabilities for a range of between-experiment variance values $b$. From Equation 1 we see that the distributional significance of a given result $p_{sig}(t|b)$ falls with falling $b$, so that if a result is significant at level $\alpha$ for some variance ratio $B$, it is significant at the same level for all $b<B$. A similar result holds for $p_{rep}(t|b)$: it can be shown that the probability of replication of a given result $p_{rep}(t|b)$ rises with falling $b$ to a maximum at some value $b_{max}\approx|d|/\sqrt{N}$, so that $p_{rep}(t|b)\geq p_{rep}(t\mid B)$ for almost all $b\leq B$ (see Appendix $1$ for proof). If we have some reasonable upper limit $B$ on the variance ratio for experiments (such that the actual unknown variance ratio for any given experiment is a priori expected to be less than this maximum), then we can define ‘generic’ estimates for distributional significance and replication as
$\hat{p}_{sig}=2T_{\nu}\left(-|t|/\sqrt{1+BN}\right)$
and
$\hat{p}_{rep}=T_{\nu_{r}}\left(\frac{|t|B\sqrt{N\,N_{r}}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)(1+bN)\sqrt{1+BN_{r}}}{\sqrt{(1+BN+BN_{r})(1+BN)}}\right)$
and if for a given result $t$ we have $\hat{p}_{sig}\leq\alpha$ and $\hat{p}_{rep}=\rho$, then our result $t$ is significant at level $\alpha$, and has a probability of replication greater than $\rho$, conditional on the assumption that between-experiment variance is less than $B$ times within-experiment variance.
What value should be chosen for this bound $B$? Differences in means across experiments can arise as a consequence of within experiment variance and as a consequence of differential factors affecting experimental means (if the means in one experiment is affected by some confounding factor not present in the other, we expect their means to differ systematically as a function of that confound). Since replication experiments are necessarily designed to be as similar as possible to each other, we expect the effect of these between-experiment confounds to be weaker than the effect of within-experiment variance (if between-experiment confounds have a strong influence on experimental means, then the experiments are not, in fact, replications: they differ in some important aspect). Given that the variance ratio $b$ represents the ratio of between-experiment to within-experiment variance, we thus expect that $b<1$ will hold for replications and set a maximum bound of $B=1$.
Note, however, that the lower the value of $B$, the lower the value of $\hat{p}_{sig}$ and the higher the value of $\hat{p}_{rep}$ (and so the more likely we are to judge a result as significant and replicable). This means that lower values of $B$ (those that may underestimate the true level of between-experiment variance) have a higher type $1$ error rate but a lower type $2$ rate, while higher values of $B$ (that may overestimate the true level of between-experiment variance) have a lower type $1$ error rate and a higher type $2$ rate. For conservative rejection of the null (only rejecting the null when the evidence is very strong), we should therefore test against a bound of $B$ close to $1$; for more liberal rejection of the null (being willing to reject the null even when the evidence is weak), we should test against a bound of $B$ close to 00.
3 Estimating significance and replication
The previous section derived distributional null expressions for significance, power and replication assuming known variance ratio $b$ or bounds on that ratio. These expressions are both computationally complex, and depend on this unknown value. In this section we derive sample estimators for $b$, and give simple and easily used closed-form estimates for significance and replication in terms of the standard $T$ distribution. We also describe extensions of this approach to different forms of $t$-test and to tests of regression, linear correlation, and contingency.
3.1 Sample Estimators for $b$
Our estimates $\hat{p}_{sig}$ and $\hat{p}_{rep}$ are in terms of generic bounds on $b$. Given replication data, however, we can estimate $b$ in terms of its component variances, producing more specific expressions for $p_{sig}$ and $p_{rep}$. Estimating $\sigma^{2}$ is straightforward: given some experiment measuring a normally distributed variable $X\sim\mathcal{N}(\mu_{i},\,\sigma_{i}^{2})$ with unknown within-experiment variance $\sigma_{i}^{2}$ and degrees of freedom $\nu$, the unbiased estimator for $\sigma^{2}$ is
$S_{i}^{2}=\frac{1}{\nu}\sum_{i=1}^{N}(X_{i}-\overline{X})^{2}$
the sample variance of $X$. We wish to give an unbiased estimator for $\sigma_{0}^{2}$, which we write as $S_{0}^{2}$. We obtain this estimator by considering the variance of the means in some set of replications of an experiment similar to ours, and subject to the same broad set of random confounding factors, both within and between experiments. Since in psychology many such factors are related to differences between individual participants (for within experiment variance) and to systematic differences between groups of participants (for between experiment variance), such estimates of the between-experiment variance of means can potentially be applied widely.
Suppose we have $K$ experiments all measuring the same variable $X$. In each experiment $i$ we assume that $X\sim\mathcal{N}(\mu_{i},\,\sigma_{i}^{2})$ for different values of $\mu_{i}$ and $\sigma_{i}^{2}$, where $\mu_{i}\sim\mathcal{N}(\mu_{0},\,\sigma_{0}^{2})$ ($\sigma_{0}^{2}$ representing the between-experiment variance of the mean). We assume that each experiment $i$ has $N_{i}$ measurements of $X$ with sample mean $\overline{X}_{i}$ and within-experiment sample variance $S_{i}^{2}$, and take $\overline{X}_{*}$ to the mean of these sample means $\overline{X}_{i}$ and
$S^{2}_{*}=\frac{1}{K-1}\sum_{i=1}^{K}(\overline{X}_{i}-\overline{X}_{*})^{2}$
to be the sample variance of these sample means. Then since each $\overline{X}_{i}$ is distributed as
$\overline{X}_{i}\sim\mathcal{N}(\mu_{i},\sigma_{0}^{2}+\sigma_{i}^{2}/N_{i})$
the theoretical variance of their overall mean $\overline{X}_{*}$ is given by
$\frac{1}{K}\sum_{i=1}^{K}(\sigma_{0}^{2}+\sigma_{i}^{2}/N_{i})=\sigma_{0}^{2}+\frac{1}{K}\sum_{i=1}^{K}\sigma_{i}^{2}/N_{i}$
and equating variances and noting that $S^{2}_{i}$ is an unbiased estimator of $\sigma^{2}_{i}$ this gives an estimate for $\sigma_{0}^{2}$ of
$S_{0}^{2}=S^{2}_{*}-\frac{1}{K}\sum_{i=1}^{K}S_{i}^{2}/N_{i}$
This is the Hedges estimator for heterogeneity in meta-analysis, and is an unbiased estimator of $\sigma_{0}^{2}$, even when $\sigma_{i}^{2}$ is unknown and is estimated by $S_{i}^{2}$ (see e.g. Hedges,, 1983; Viechtbauer,, 2005; Veroniki et al.,, 2016). Since the within-experiment sample variance of the mean $S_{i}^{2}/N_{i}$ declines to 00 with increasing $N_{i}$, the expression $(K-1)S_{0}^{2}/\sigma_{0}^{2}$ approximately follows a $\chi^{2}$ distribution with degrees of freedom $\nu_{0}=K-1$ (with the error of the approximation falling with increasing $N_{i}$).
Note that in the case where we have only seen $1$ experiment with mean $\overline{X}$ and sample variance $S^{2}$ and no data on the variation of sample means across experiments, the above expression for $S_{*}^{2}$ is undefined and so an estimate for $S_{0}^{2}$ cannot be given. Similarly, since estimates for $\sigma_{0}^{2}$ necessarily depend on accurate estimates of the component variances $\sigma_{i}^{2}$, when $S^{2}_{0}<0$ we can conclude that some sample sizes $N_{i}$ are too small to give the required accuracy; and again, an estimate for $S_{0}^{2}$ cannot be given.
Given unbiased estimators $S^{2}$ and $S_{0}^{2}$ we take $\hat{b}=S_{0}^{2}/S^{2}$ as our estimate for the ratio $b=\sigma^{2}_{0}/\sigma^{2}$, and because both $S^{2}$ and $S_{0}^{2}$ approximately follow $\chi^{2}$ distributions, we see that $b$ approximately follows an $F_{\nu,\nu_{0}}$ distribution scaled by this estimate $\hat{b}$ (with the error in the approximation falling with $\nu_{0}$). The variable $p_{sig}$ is then distributed as a function of the distribution of $b$, and the expected value of that distribution given $\hat{b}$ is
$p_{sig}(t|\hat{b})=\int_{0}^{\infty}2T_{\nu}\left(\frac{-|t|}{\sqrt{1+b\hat{b}N}}\right)f_{\nu,\nu_{0}}(b)\mathrm{d}b$
where $f_{\nu,\nu_{0}}$ is the probability density of that $F$ distribution. This is our theoretical model for distributional significance, and when $p_{sig}(t|\hat{b})<\alpha$ we can say that our result $t$ gives statistically significant evidence against the null hypothesis, given the estimated within-experiment and between-experiment variances.
Assuming that sample variances $S^{2}$ and $S_{r}^{2}$ in the two experiments are both estimates of the same underlying within-experiment variance $\sigma^{2}$, then since these two sample variances both follow a $\chi^{2}$ distribution with the same mean and with degrees of freedom $\nu$ and $\nu_{r}$, their ratio $c$ follows an $F$ distribution $F_{\nu,\nu_{r}}$. Similarly taking $\hat{b}=S_{0}^{2}/S^{2}$ as our estimate for the ratio $b$ we see that the variable $p_{rep}$ is distributed as a function of the joint distribution of $b$ and $c$. Since we don’t know the sample variance $S_{r}^{2}$ (our replication experiment has not yet been carried out) we simply set $\hat{c}=1$ and the expected value for $p_{rep}$ given estimate $\hat{b}$ is
$p_{rep}(t|\hat{b})=\int_{0}^{\infty}\int_{0}^{\infty}T_{\nu_{r}}\left(\frac{|t|\sqrt{N\,N_{r}}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)(1+b\hat{b}N)\sqrt{1+b\hat{b}N_{r}}}{\sqrt{(1+b\hat{b}N+b\hat{b}N_{r})(1+b\hat{b}N)}}\right)f_{\nu,\nu_{0}}(b)f_{\nu,\nu_{r}}(c)\,\mathrm{d}b\,\mathrm{d}c$ (5)
This is our theoretical model of the probability of result $t$ being replicated in an experiment with $N_{r}$ measurements, given an estimate of between-experiment variance in the means of a set of similar experiments and assuming the sample variance in that replication is close to $S^{2}$. Note that this estimate for replication is conditional on evidence against that null hypothesis from the experimental result $t$. This means that if result $t$ does not give evidence against the null hypothesis (that is, if $p_{sig}$ is high) then we are not justified in ‘updating away’ from that hypothesis, and our estimate of $p_{rep}$ does not apply. In other words, we only have evidence for replication when we have evidence against the null.
These integral and double-integral estimates for $p_{sig}$ and $p_{rep}$ are, however, somewhat computationally expensive to calculate in practice; and further, these calculations can be ‘fragile’: dependent in opaque ways on the numerical integration algorithm used, especially for values of $\hat{b}$ close to 00. In the next section we give more practically useful estimates for $p_{sig}$ and $p_{rep}$, expressed in terms of the standard $T$ distribution.
3.2 Closed-form significance and replication estimates
The difference between point-form and distributional significance testing is most important for large sample size $N$, because in that situation between-experiment variance $\sigma_{0}^{2}$ dominates the sample variance of the mean $\sigma^{2}/N$. Here we derive approximate estimates for significance and replication for large values of $N$: specifically, values such that $S^{2}$ is approximately equal to $\sigma^{2}$ and such that $\sigma/\sqrt{N}$ is much smaller than $\sigma_{0}$, which means that
$\frac{bN}{1+bN}\approx 1$
Then we have
$p_{sig}(t|b)=2T_{\nu}\left(\frac{-|t|}{\sqrt{1+bN}}\right)\approx 2\Phi\left(\frac{-|t|}{\sqrt{1+bN}}\right)\approx 2\Phi\left(\frac{-|\overline{X}|}{(\sigma/\sqrt{N})\sqrt{\frac{\sigma_{0}^{2}}{\sigma^{2}}N}}\right)=2\Phi\left(\frac{-|\overline{X}|}{\sqrt{\sigma_{0}^{2}}}\right)$
Since by assumption $\sigma/\sqrt{N}$ is much smaller than $\sigma_{0}$, we see that $\overline{X}$ approximately follows the distribution $\mathcal{N}(0,\sigma_{0}^{2})$. Since $S_{0}^{2}$ is an unbiased sample estimator for $\sigma_{0}^{2}$ such that $\nu_{0}S_{0}^{2}/\sigma_{0}^{2}$ approximately follows the distribution $\chi^{2}_{\nu_{0}}$, this means that
$\Phi\left(\frac{-|\overline{X}|}{\sqrt{\sigma_{0}^{2}}}\right)\approx T_{\nu_{0}}\left(\frac{-|\overline{X}|}{\sqrt{S_{0}^{2}}}\right)=T_{\nu_{0}}\left(\frac{t}{\sqrt{\hat{b}N}}\right)$
(Note the change in degrees of freedom for the $T$ distribution, from $\nu$ to $\nu_{0}$), and we define our general expression for distributional significance as
$p_{sig}=2T_{\nu_{0}}\left(\frac{-|t|}{\sqrt{\hat{b}N}}\right)$
For replication we have
$\begin{split}p_{rep}(t|b)&=T_{\nu_{r}}\left(\frac{|t|b\sqrt{N\,N_{r}}}{\sqrt{(1+bN)(1+b(N+N_{r}))}}-\frac{T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)\sqrt{(1+bN)(1+bN_{r})}}{\sqrt{1+b(N+N_{r})}}\right)\\ \end{split}$
and since by assumption $\sqrt{1+bN}\approx\sqrt{bN}$, we have the approximation
$\begin{split}p_{rep}(t|b)&\approx T_{\nu_{r}}\left(\sqrt{\frac{N_{r}}{N+N_{r}}}\left[|t|-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)\sqrt{1+bN}\right]\right)\\ \end{split}$
By assumption we also have $S^{2}\approx\sigma^{2}$ so that $b=\sigma_{0}^{2}/\sigma^{2}\approx\sigma_{0}^{2}/S^{2}$ and so we can rearrange to get
$p_{rep}(t|b)\approx T_{\nu_{r}}\left(\sqrt{\frac{\hat{b}NN_{r}}{N+N_{r}}}\left[\frac{|t|}{\sqrt{\hat{b}N}}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)\sqrt{\frac{1}{\hat{b}N}+\frac{\sigma_{0}^{2}}{S_{0}^{2}}}\right]\right)$
Since $\sigma_{0}^{2}/(\nu_{0}\,S_{0}^{2})$ approximately follows the inverse distribution $\chi^{-2}_{\nu_{0}}$, and this distribution has a mean of $1/(\nu_{0}-2)$ for $\nu_{0}>2$, this gives an expected replication probability of
$p_{rep}=T_{\nu_{r}}\left(\sqrt{\frac{\hat{b}NN_{r}}{N+N_{r}}}\left[\frac{|t|}{\sqrt{\hat{b}N}}-T^{-1}_{\nu_{r}}\left(1-\alpha/2\right)\sqrt{\frac{1}{\hat{b}N}+\frac{\nu_{0}}{\nu_{0}-2}}\right]\right)$
which defines our general expression for replication.
These $p_{sig}$ and $p_{rep}$ expressions are closed-form estimates for distributional significance and replication, analogous to the point-form $t$ test for significance in that they use the $\chi^{2}$ distribution to replace the unknown variance $\sigma^{2}_{0}$ with its estimate $S_{0}^{2}$. We expect these expressions to be less accurate in terms of estimating significance and replication probabilities than the full integral expression given in Equation 5 (with the difference falling as sample size $N$ rises) but these closed-form estimates have the advantages of simplicity, computational speed, and robustness.
These expressions also clarify the general relationship between distributional significance and replication: given a result $t$ which exactly meets the criterion for significance we have
$|t|=T^{-1}_{\nu_{0}}\left(1-\alpha/2\right)\sqrt{\hat{b}N}$
and so substituting the probability of replication for this result is
$p_{rep}(t|b)=T_{\nu_{r}}\left(T^{-1}_{\nu_{0}}\left(1-\alpha/2\right)\,\sqrt{\frac{\hat{b}NN_{r}}{N+N_{r}}}\left[1-\sqrt{\frac{1}{\hat{b}N}+\frac{\nu_{0}}{\nu_{0}-2}}\right]\right)$
The innermost term is negative and approaches 00 with rising $\nu_{0}$ and $\hat{b}N$, and so $p_{rep}$ for such a result is less than, but approaching, $0.5$. This means that non-significant results necessarily have $p_{rep}<0.5$ in this model and significant results have $p_{rep}\geq 0.5$ (except in limiting cases where $N$ is small and $p_{sig}$ is close to $\alpha$). In other words, where the standard point-form significance measure $p$ tells us nothing about the probability of replication of a given result, the distributional significance measure $p_{sig}$ gives direct information about the probability of replication: if $p_{sig}<\alpha$ for a given result then we expect that result to replicate at least $50\%$ of the time (with the probability of replication increasing as $p_{sig}$ falls).
3.3 Extension to related statistical tests
We’ve presented distributional significance testing here in terms of a generalised $t$-test, and so this approach applies to each distinct form of that test. To apply this model to a one-sample $t$-test on a variable $X$ with sample size $N$, we set $\nu=N-1$ and take $S^{2}$ to be the sample variance of that variable. To apply to a paired $t$-test on two variables (assuming $N$ pairs) we take the $X$ to represent the difference in pairs and $\overline{X}$ and $S^{2}$ to be the average difference between pairs and the variance of that difference, and $\nu=N-1$. To apply to an unpaired or independent $t$-test on two variables (with $N$ measurements in total), we take $\overline{X}$ to be the difference between the means of each variable, $N$ to be the sample size, and $S^{2}$ to be the pooled sample variance with degrees of freedom $\nu=N-2$. (Note that we implicitly assume an equal number of samples for the two variables in an unpaired test. This is because the probability of replication of a given result in a future experiment on two unpaired variables will depend on the sample sizes for those two variables in that experiment, which we cannot know. Assuming that these sample sizes vary randomly the expected sample size for each variable will be $N/2$; and so we use the calculated replication probability assuming equal sample sizes as our estimate of the probability of replication.)
The approach also applies to tests of linear regression, correlation and contingency, which can all be related to the $t$-test. Applying the distributional approach to tests of independence for correlation and linear regression gives $p_{sig}$ and $p_{rep}$ expressions as above, albeit calculated with sample slope $M$ replacing sample mean $\overline{X}$, predictor sum-of-squares (which we write as $Q$) replacing sample size $N$, and mean squared error of the response estimate replacing sample variance. This is because, with these replacements, such tests of independence are simply $t$-tests (see Appendix $2$ for a derivation of these results). Note that in these correlation tests the sum of squares of the predictor variable is the continuous-valued analog of the discrete sample size $N$ in a $t$-test, and is treated as a controlled variable: where sample size gives the number of discrete measurements made in a $t$-test, predictor sum-of-squares gives the range of continuous measurements in a test of correlation. This means that the significance and replication probability of these tests depends on the choice of predictor and response variables: an estimate for the probability of replication of a given result in an experiment where $X$ is the predictor (and so assumed to be controlled in replications) will necessarily be different from the probability of replication of that result when $Y$ is the predictor and the $Y$ sum of squares is assumed to be controlled in replications (because the two replication experiments themselves are different).
Code to calculate all these tests of distributional significance and replication is available online at https://osf.io/ep86y. This code also tests the distributional model’s approach to significance and replication against the Many Labs 1 replication dataset, as described in the next section.
3.4 Comparing with point-form approaches to replication
How does this distributional model compare with other approaches to replication? The most influential method for estimating the probability of replication $p_{rep}$ is that of Killeen (Killeen, 2005a, ; Killeen, 2005b, ; Killeen,, 2007; Sanabria and Killeen,, 2007), which for a number of years was recommended by the journal Psychological Science as an alternative to standard hypothesis tests. This approach does not estimate the probability of obtaining a statistically significant result in a second experiment: instead it estimates the probability of obtaining a result in a second experiment that is the same sign as the original result. In a one-sample test with sample size $N$ and effect $d=\overline{X}/S$, Killeen’s $p_{rep}$ calculates the probability of obtaining a result of the same sign as $d$ by estimating the variance of the original effect as $\sigma^{2}_{d}\sim N/(N-4)$, taking the variance in the ensemble of original and replication to be $2\sigma^{2}_{d}$, and assuming that effects are normally distributed, so that Killeen’s $p_{rep}$ for a given $d$ is approximately
$\Phi\left(\frac{|d|}{\sqrt{2}\sigma_{d}}\right)=\Phi\left(\frac{|\overline{X}|}{\sqrt{2}S}\left[1-\frac{4}{N}\right]\right)$
Researchers have pointed out a number of fundamental problems with this approach to replication: that there are errors in the derivation of Killeen’s $p_{rep}$ (Doros and Geier,, 2005); that Killeen’s $p_{rep}$ is simply a function of the point-form $p$ and so gives no further information (Maraun and Gabriel,, 2010); that Killeen’s $p_{rep}$ estimates an aggregated probability over all results and experiments, rather than the probability of a given result being replicated in a particular experiment (Miller,, 2009; Miller and Schwarz,, 2011); that replication probabilities must be estimated based on a prior distribution (Macdonald,, 2005); that Killeen’s $p_{rep}$ assumes all effect sizes are equally likely and so assumes a priori that the null hypothesis is false (Iverson et al., 2010b, ); and finally that the claims associated with Killeen’s $p_{rep}$ are easily misinterpreted (Iverson et al., 2009a, ; Iverson et al., 2009b, ; Iverson et al., 2010a, ).
Our distributional null hypothesis expression for $p_{rep}$ resolves the problems in Killeen’s approach. Our $p_{rep}$ is not simply a function of the point-form $p$ and is an estimate for replication in a particular experiment (in that it depends on the ratio of between-experiment to within experiment variance $b$, the ratio of within-experiment variance in the original experiment and the replication $c$, and the sample sizes of the original and replication experiments). Our $p_{rep}$ is specifically based on a distribution for the null hypothesis, does not assume all effect sizes are equally likely (the larger an effect size the less likely it is under this distribution), and does not assume the null hypothesis is false. Finally, since our result is based on well-known properties of the Normal and $t$ distributions, errors in the derivation are (hopefully) unlikely, and the associated claims have a clear interpretation.
4 Predicting Replication in the Many Labs 1 Dataset
We test this distributional approach using data from the first ‘Many Labs’ replication project (Klein et al.,, 2014). This involved the replication of $16$ different experiments investigating a variety of classic and contemporary psychological effects. Each experiment was originally published in the cognitive or social psychology literature, and was replicated by researchers in around $36$ different sites ($25$ in the US, $11$ international): each individual instance of one of these tasks thus had around $35$ replications with the same design but with varying sample size and sample variance values (some tasks had less than $36$ experiments; there were $574$ experiments in total, with $400$ occurring at US sites). We chose to use this dataset to test the distributional null approach because the raw response data for all experiments at all sites is available, allowing us to compute the various estimators required in the distributional model, and because the large number of replications of each individual task allows an accurate estimate of the actual replication rate for these experiments, which we can use to test the replication probability predicted for those experiments by our model. Note that this dataset does not include published results from any of the original experiments being tested, but only results from planned replications across the $36$ sites (so minimising the impact of publication bias).
These experiments covered a wide range of topics, including the impact of question phrasing on whether speeches for or against democracy should be allowed/forbidden; the effect of ‘anchoring’ on estimates of the distance from San Francisco to NYC, the number of babies born per year in the US, the population of Chicago, or the height of Mount Everest; agreement with quotes depending on they are attributed to George Washington/Osama Bin Laden; the effect of priming with the US flag/currency on political opinions; and the effect of imagined contact with Muslims on reductions in prejudice. A number of these topics were US-specific, and so experimental “replications” of these results in non-US sites are subject to clear confounds (we would expect there to be differences between US and non-US sites in knowledge about US distances and US city populations, and in attitudes towards George Washington and the US flag or currency, for example [^2] Such differences are evident even in knowledge about the height of Mount Everest: on average fewer than $2\%$ of estimates at US sites fell within $1\%$ of the true height, with all US sites having less than $6\%$ of estimates in that range. Of the $11$ international sites, however, one had $35\%$ of estimates within $1\%$ of the true height, one $26\%$ and one $20\%$, and on average more than $10\%$ of international estimates were within $1\%$ of the true height.). For this reason we limit our analysis to results from US sites only.
4.1 Methods
All data, analysis and visualisation code is publicly available on the OSF repository (Costello,, 2022). Data were analyzed using R, version $4.0.5$ (R Core Team,, 2021) and the packages ggplot2 version $3.3.5$ (Wickham,, 2016) and xtable $1.8.4$(Dahl et al.,, 2019). The R analysis script downloads the Many Labs $1$ dataset from the project’s OSF repository and carries out all analyses described below.
Experiments in the Many Labs dataset varied widely in significance, with some having significant results in all or almost all replications, and others having only a few significant results. In our analysis we use the independent and dependent measures from the Many Labs 1 tasks (and the corresponding criteria for inclusion/exclusion of responses as flagged in the dataset); Table 1 gives summary information on these tasks for US sites.
Table 1: Summary information for US sites in Many Labs 1 replication project, with $25$ individual experiments for each experimental task. The table also gives the estimated sample variance of the means across experiments in each task ($S^{2}_{0}$), the proportion of experiments in each task that were statistically significant in two-sided point-form and distributional tests ($p$ or $p_{sig}$ less than $0.05$) and where the effect was in the same direction as in the original experiment (this is equivalent to significance at $0.025$ in a one-sided point-form or distributional test with the direction of the test corresponding to that of the original effect). The $p$ and $p_{sig}$ values of the aggregated dataset for each experimental task are also shown (with between-experiment variance for a given task taken to be $S^{2}_{0}$ and within-experiment variance calculated from the aggregated dataset for that task).
significant at $\alpha=0.05$ and in
Task
Total test $S_{0}^{2}$ $\hat{b}$ same direction as original
participants proportion    aggregated
(mean) $p$ $p_{sig}$ $p$ $p_{sig}$
1
Allowed/forbidden
4934 $\chi^{2}$ 0.005 0.041 1.00 1.00 $10^{-325}$ $10^{-10}$
2
Anchoring - babies born
4481 t $6.4\times 10^{6}$ 0.06 1.00 1.00 $10^{-325}$ $10^{-4}$
3
Anchoring - Mt. Everest
4575 t $9.3\times 10^{6}$ 0.11 1.00 1.00 $10^{-325}$ 0.0004
4
Anchoring - Chicago
4302 t $8.1\times 10^{10}$ 0.077 1.00 1.00 $10^{-325}$ 0.001
5
Anchoring - distance to NYC
4422 t $5.0\times 10^{4}$ 0.05 1.00 0.76 $10^{-266}$ 0.01
6
Relation between Implicit and Explicit math attitudes
4383 r 0.080 0.12 0.88 0.52 $10^{-148}$ 0.01
7
Retrospective gambler fallacy
4690 t 0.640 0.12 0.88 0.00 $10^{-98}$ 0.15
8
Low vs High category scales
4664 $\chi^{2}$ 0.004 0.077 0.80 0.28 $10^{-57}$ 0.01
9
Gain vs. loss framing
4925 $\chi^{2}$ 0.007 0.029 0.84 0.68 $10^{-91}$ 0.001
10
Sex differences in implicit math attitudes
4558 t 0.010 0.05 0.76 0.00 $10^{-56}$ 0.15
11
quote attribution
4964 t 0.260 0.063 0.56 0.00 $10^{-36}$ 0.23
12
sunk costs
4967 t 0.110 0.032 0.36 0.00 $10^{-19}$ 0.19
13
norm of reciprocity
4928 $\chi^{2}$ 0.008 0.04 0.44 0.20 $10^{-28}$ 0.07
14
imagined contact
4973 t 0.110 0.034 0.16 0.00 $10^{-7}$ 0.34
15
Flag priming
4896 t 0.012 0.017 0.00 0.00 0.20 0.45
16
Currency priming
4970 t 0.016 0.021 0.00 0.00 0.40 0.49
For every individual experiment, in this script we calculated the within-experiment sample variance $S^{2}$; for each experimental task we calculated the between-experiment sample variance $S_{0}^{2}$ (in terms of sample means $\overline{X}$ or slopes $\hat{M}$). To assess the distributional model of significance against these experimental results, we used these sample variances to calculate the distributional significance estimate $p_{sig}$ for each experiment, and compared against the point-form significance $p$ for each experiment.
Figure 1: (Left) Histogram of the standardised sample means, $z=(\overline{X}_{i}-\overline{X}_{*})/S_{0}$, for US sites for each task in the Many Labs 1 replication dataset, in bins of size $0.5$ with colour representing task. The line shows the standard Normal distribution $\mathcal{N}\left(0,1\right)$, scaled by bin size $\times$ number of observations to match the histogram scale. (Right) QQ plot showing quantiles of the standardised sample means, $z=(\overline{X}_{i}-\overline{X}_{*})/S_{0}$, for US sites for each task against quantiles of the standard Normal distribution.
To assess the distributional model of replication, we grouped all experiments for a given task into ‘predictor-target’ pairs and used the results of the predictor experiment to calculate the probability of replication $p_{rep}$ for that particular target experiment (and a given value $\alpha$). We then compared the predicted replication probability for each target experiment against the observed rate of successful replication in those target experiments.
4.2 Distribution of means across experiments
To test the assumption that experimental means vary between experiments following a Normal distribution, for each task we calculate $\overline{X}_{*}$ (the average of sample means $\overline{X}_{i}$ for each replication of that task; for correlation tests we take $\overline{X}_{i}=\hat{M}_{i}$) and $S^{2}_{0}$ (the sample estimate of the variance of means or slopes $\mu_{i}$ around the overall mean $\mu_{0}$), and normalise to $z=(\overline{X}-\overline{X}_{*})/S_{0}$.
Figure 1(left) plots a histogram of these $z$ values and shows consistent variance in experimental means $\overline{X}$ for all tasks, with $z$ scores for means following the standard Normal distribution. Figure 1(right) plots the quantiles of these $z$ values for each individual task against quantiles of the standard Normal distribution, confirming that individual experimental means for each task follow this distribution. Shapiro-Wilk tests showed no significant evidence against the Normal distribution for all but $1$ task (point-form $p>0.05$ for all but $1$ task; note that since the specific model here is one where experimental means $\mu_{i}$ are normally distributed around a point-form mean $\mu_{0}$, the point-form test for normality is appropriate in this case). Together these results support the distributional model’s assumption that means of experimental replications vary randomly following an approximately Normal distribution.
4.3 Comparing point-form and distributional significance
Statistical significance in this dataset was originally assessed in a two-sided test relative to the point-form $p$ at $\alpha=0.05$, with a given experimental result counted as significant when it was both significant in this two-sided test and where the effect was the same direction as that seen in the originally published experiment (equivalent to a one-sided test in the original direction with $\alpha=0.025$). We follow this definition for distributional significance.
As an initial check we used numerical integration methods to calculate, for each experiment, the theoretical value $p_{sig}(t|\hat{b})$ given that experiment’s $t$ value and the $\hat{b}$ estimate for that task, and compared against the closed-form expression $p_{sig}$ for that experiment. $p_{sig}p(t|\hat{b})$ and $p_{sig}$ values were almost identical (correlation $r=0.99$), supporting the closed-form approximations, and so we used only the closed-form approximation $p_{sig}$ in our analysis.
To compare point-form and distributional measures of significance ($p$ and $p_{sig}$) we calculate, for each task, the $p$ value for results from each US site and record the percentage of sites which have $p\leq 0.05$. Taking $\hat{b}=S_{0}^{2}/S^{2}$ for each task and experiment we similarly calculate $p_{sig}$ values for each site and record the percentage of sites with $p_{sig}\leq 0.05$. Table $1$ compares, for US sites, the percentage of successful replications of a given task at the $\alpha=0.05$ level under point-form ($p$) and distributional ($p_{sig}$) tests, and also reports the $p$ and $p_{sig}$ values for the aggregated data (across all US sites) for each task. Point-form $p$ values for aggregated data are unreasonably extreme: $4$ tasks had $p<10^{-325}$ (the computational minimum); a task with only $16\%$ successful replications relative to the point-form $p$ had aggregate significance of $p<10^{-7}$. Distributional significance values $p_{sig}$ were orders of magnitude more conservative, both in terms of proportion of successful replications and in the aggregate measure.
Figure 2: (Left) Scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ point-form $p$ for all experiments in all tasks in the Many Labs 1 replication dataset for US sites (note the difference in scales). Lines indicate statistical significance at levels $\alpha=0.05$ (dotted), $\alpha=0.025$ (dashed) and $\alpha=0.01$ (solid). (Right) Analogous scatterplot of $\log_{10}$ distributional $p_{sig}$ against $\log_{10}$ Bayes Factor $BF_{10}$. Horizontal lines indicate Bayes Factor levels of $3$ (dotted: $BF_{10}\geq 3$ being typically interpreted as moderate evidence against the null), $10$ (dashed: strong evidence against the null) and $30$ (solid:extremely strong evidence against the null). Points for $9$ experiments, all with $p<10^{-40}$ ($BF_{10}>10^{40}$) and ranging as low as $p<10^{-325}$ ($BF_{10}>10^{325}$), are not shown.
We also compared $p_{sig}$ and point-form $p$ values for each individual result on a $\log_{10}$ scale (see Figure 2(Left)). There are many extremely low $p$ values, with $37\%$ of experiments having $p<10^{-5}$ and $22\%$ of experiments having $p<10^{-15}$. Distributional $p_{sig}$ values were again orders of magnitude more conservative ($94\%$ greater than $10^{-5}$ and all greater than $10^{-10}$). Many results were significant under point-form but not distributional tests: $40\%$ of results with $p<0.05$ had $p_{sig}>0.05$ and $11\%$ of results with $p<10^{-5}$ had $p_{sig}>0.05$. Even apparently highly significant results (in point-form tests) can turn out to be consistent with the null hypothesis when between-experiment variation is taken into account.
4.4 Comparing predicted and observed replication probabilities
To assess the distributional model of replication, for each of the $16$ tasks we generated every possible pair of experiments from two different US sites and calculated the replication probability $p_{rep}$ for the second or target experiment in each pair based on the result $t$ in the first or predictor experiment. This calculation used the between-experiment variance of means for the task overall, $S_{0}^{2}$, the within-experiment variance in the predictor experiment $S^{2}$ and the sample size of the predictor experiment $N$, and, given these values, generating the probability of replication for an experiment with the sample size $N_{r}$ of the paired target experiment. This calculation assumed $c=1$ and so did not make use of the sample variance of the target experiment (and so represents the calculation a researcher would carry out when they have observed a surprising result in an experiment with sample size $N$ and within-experiment variance $S^{2}$, and are estimating the probability of replicating that result in a planned replication with sample size $N_{r}$ but unknown within-experiment variance). We calculated this replication probability for a range of commonly used significance levels $\alpha=,0.05,0.025,0.01,0.005,0.001$. For each target experiment and each level $\alpha$ we compared this predicted replication probability against a dichotomous successful replication variable ($p_{sig}\leq\alpha$) which indicated whether the target experiment did, in fact, replicate successfully at level $\alpha$.
As an an initial check we calculated, for each pair of experiments in a given task, both the closed-form approximation $p_{rep}$ and the double-integral expression $p_{rep}(t|\hat{b})$. As before, these were almost identical (correlation $r=0.94$) supporting the closed-form approximations, and so we primarily make use of the closed-form approximation $p_{sig}$ in our analysis (Figure $3$ gives a visual comparison of $p_{rep}$ and $p_{rep}(t|\hat{b})$ values).
Table 2: Correlation between predicted and observed replication rates at each level of $\alpha$, with distributional significance $\hat{p}_{sig}(B=1)$ and replication probability $\hat{p}_{rep}(B=1,\alpha=0.05)$ for these correlations. The table also gives the relevant parameters characterising each correlation: slope, predictor sum-of-squares $Q$ and response variation or error $S^{2}$.
[…]
($p_{sig}<\alpha$) $p_{rep}$ vs ($p_{sig}<\alpha$) for $r$ for $r$
0.05 9600 48 0.76 0.002 1 0.95 1534 0.11
0.025 9600 40 0.78 0.0007 1 0.97 1499.8 0.09
0.01 9600 34 0.80 0.0002 1 1.01 1367.5 0.08
0.005 9600 30 0.82 0.00003 1 1.05 1224 0.07
0.001 9600 20 0.78 0.00002 1 1.04 846.18 0.06
all 48000 34 0.80 0.0002 1 1.01 6738.7 0.083
Correlation between $p_{rep}$ values and dichotomous successful replication values ($p_{sig}\leq\alpha$) was $r=0.80$ across the entire set with a slope of $1.01$. Point form $p$ for these correlations between predicted and observed replication was less than $10^{-325}$ (the computational minimum) in all cases. Table $2$ gives the correlation, significance and replication values for each level of $\alpha$ and overall. The table also gives the necessary statistics for future calculations of between-experiment variance using these results (the sample slope, sum of squares of the predictor variable, and sample variance or error of the response variable).
Figure 3: Predicted vs observed replication rate (for $\alpha=0.05,0.025,0.01,0.005,0.001$) for target experiments in the Many Labs 1 replication dataset for US sites, grouped by $p_{rep}$ in bins of size $1/40$. Bubble size represents the number of predictor-target pairs in each bin (sizes range from $40$ pairs for the smallest to $2800$ for the largest bubble). Hollow bubbles contain only non-significant and solid bubbles only significant predictor experiments. The left graph shows results for the double-integral $p_{rep}(t|\hat{b})$ replication estimate, the right for the closed-form $p_{rep}$ estimate. Correlation between predicted and observed replication rates for bubbles was high for $p_{rep}(t|\hat{b})$ ($r=0.91,\hat{p}_{sig}<10^{-11}$) and for the approximation $p_{rep}$ ($r=0.81,\hat{p}_{sig}<10^{-7}$, $\hat{p}_{rep}=1.0$ with $\alpha=0.05$ and $B=1$ for both expressions).
To illustrate the relationship visually, we combined the sets of predictor-target pairs for these values of $\alpha$ into a single collection, and grouped those pairs into bins of size $1/40$ by level of $p_{rep}$ and task: each bin contained a set of predictor-target pairs involving the same task with approximately the same replication probability $p_{rep}$ produced from the predictor experiment for the target (for the value $\alpha$ initially applied to that pair). To distinguish between predictions from statistically significant and non-significant predictor experiments, we carried out this grouping process once for all pairs for which the predictor experiment was significant at $p_{sig}\leq\alpha$, and a second time for all pairs where the predictor experiment was non-significant. To estimate the true rate of successful replication for target experiments in each bin we calculated the proportion of pairs in each bin for which the target experiment was significant ($p_{sig}\leq\alpha$ held for the pair’s assigned value of $\alpha$). We excluded bins containing less than $40$ predictor-target pairs, to allow a relatively accurate estimate of the actual replication rate for experiments in each bin (and to produce a matched level of resolution for predicted and observed replication rates). Figure 3 shows a bubbleplot of predicted and observed replication rates for the double integral $p_{rep}(t|b)$ and for $p_{rep}$. As the figure shows there was a reliable linear relationship between observed and predicted replication rates across bins.
One concern is that the observed relationship between predicted and observed replication probabilities could simply be an artefact of the differences in significance rates across different tasks. To see the problem, assume that a statistically significant experiment necessarily produces a high predicted replication probability, while a non-significant experiment necessarily produces a low predicted replication probability. In this situation a task where all or almost all experiments were statistically significant would produce both a high predicted replication rate and (since every experiment is significant) a high observed replication rate (for the same reason), while a task where all or almost all experiments were non-significant would produce both a low predicted replication rate and a low observed replication rate. Figure 3 shows that this artefact does not explain these results: predicted and observed replication probabilities were related at the centre of the probability scale, as well as at the boundaries.
Figure 4: Sensitivity analysis. Each graph shows the predicted vs observed replication rate (for $\alpha=0.05,0.0.25,0.01,0.005,0.001$) for target experiments in the Many Labs 1 replication dataset for US sites as in Figure 3, but with between-experiment variation $S_{0}^{2}$ multiplied by scale $e$.
4.5 Sensitivity to estimates of between-experiment variance
The predicted probability of replication in the distributional model, $p_{rep}$, depends on an estimate of between-experiment variance. In this section we ask how sensitive the accuracy of this predictive probability is to the between-variance estimate used in making the prediction. To test this we re-ran our replication assessment on the US sites in the Many Labs $1$ dataset just as above, but with the between-experiment variance estimate $\hat{\sigma_{0}^{2}}$ for each task multiplied by a scaling value $e$. In the first sensitivity analysis we set $e=0.5$ for all tasks (simulating a situation where the between-experiment variance used in calculating $p_{rep}$ is underestimated by $50\%$ relative to the actual experimental value for each task); in subsequent analyses we set $e$ to $0.75$, $1.25$ and $1.5$ (calculating $p_{rep}$ with between-experiment variance estimates of $75\%$, $125\%$ or $150\%$ of the actual value).
Correlations between $p_{rep}$ values and dichotomous successful replication values for target experiments, combined across all $\alpha$ levels, were $r=0.75$ or higher for all values of $e$, indicating that even with significant over- or underestimation of between-experiment variance, $p_{rep}$ reliably predicts replication rates. Figure 4 shows the predicted/observed replication rates produced in these analyses; correlation between predicted and observed replication rates for bubbles was above $r=0.75$ for all values of $e$ ($\hat{p}_{sig}<10^{-5},\hat{p}_{rep}=1$, with $\alpha=0.05$ and $B=1$). As the figure demonstrates, replication probability tended to be underestimated when between-experiment variance was underestimated ($e<1$) but overestimated when between-experiment variance was overestimated ($e>1$), with the degree of under/overestimation in replication increasing with the degree of under/overestimation in between-experiment variance. This pattern follows from the relationship between type $1$ and $2$ error rates and the value of $b$. There was relatively little difference in prediction accuracy for $e=0.75$ and $e=1.25$, suggesting that the $p_{rep}$ is not overly sensitive to error in estimates of between-experiment variance.
5 Comparing with Bayesian approaches
This distributional model mixes ‘frequentist-style’ hypothesis testing with ‘Bayesian-style’ updating and posterior prediction for estimating replication probabilities. Advocates of the Bayesian approach may consider that, in the words of one reviewer, we are ‘teetering on the edge of Bayes already’ and ask why we do not ‘go all the way’, by abandoning tests of statistical significance and moving to alternative Bayesian approaches. In this section we address this possibility by comparing the distributional model against the best-known and most commonly used Bayesian alternative to significance testing, the Bayesian or JZS $t$-test (Rouder et al.,, 2009).
5.1 JZS Bayesian $t$-test
Given some observed data $\overline{X}$ on a parameter of interest the Bayes Factor $BF_{10}$ comparing two hypotheses $H_{1}$ and $H_{0}$ is the ratio of the marginal likelihood of $\overline{X}$ given $H_{1}$ to the marginal likelihood of $\overline{X}$ given $H_{0}$:
$BF_{10}=\frac{P(\overline{X}|H_{1})}{P(\overline{X}|H_{0})}$
Where a (point-form or distributional) $p$-value only gives evidence against the null hypothesis $H_{0}$, this measure $BF_{10}$ is taken to give evidence against $H_{0}$ and in favour of $H_{1}$: the higher $BF_{10}$, the more $\overline{X}$ gives Bayesian evidence in favour of $H_{1}$, while the higher its inverse $BF_{01}=1/BF_{10}$, the more $\overline{X}$ gives evidence in favour of $H_{0}$.
To use a Bayes Factor approach to assess evidence for or against a null hypothesis, the JZS Bayesian $t$-test (Rouder et al.,, 2009) takes the parameter of interest to be the sample mean $\overline{X}$ and takes $H_{0}$ to be the standard point-form null
$H_{0}:\overline{X}\sim\mathcal{N}(0,\sigma^{2}/N)$
with unknown variance $\sigma$, so the probability of $\overline{X}$ given $H_{0}$ is the density of this distribution at $\overline{X}$,
$p(\overline{X}|H_{0})=\int_{0}^{\infty}\mathcal{N}(\overline{X}|0,\sigma^{2})f_{\sigma}(\sigma)\mathrm{d}\sigma$
where $f_{\sigma}$ is the density of the assumed prior for $\sigma$ (which in the JZS test is the Jeffreys prior). The alternative hypothesis $H_{1}$ is that the unknown mean $\mu$ itself follows a Normal distribution $\mu\sim\mathcal{N}(0,\sigma_{\mu}^{2})$, so that the parameter of interest follows the distribution
$H_{1}:\overline{X}\sim\mathcal{N}(0,\sigma^{2}/N+\sigma_{\mu}^{2})$
(identical to our distributional null model) so that
$p(\overline{X}|H_{1})=\int_{0}^{\infty}\int_{0}^{\infty}\mathcal{N}(\overline{X}|0,\sigma^{2}/N+\sigma_{\mu}^{2})f_{\sigma}(\sigma)f_{\sigma_{\mu}}(\sigma_{\mu})\,\mathrm{d}\sigma\,\mathrm{d}\sigma_{\mu}$
where $f_{\sigma_{\mu}}$ is the prior density of $\sigma_{\mu}$. [^3] Note that this distribution is typically presented in term of a reparameterisation to effect size $\delta$ with unknown variance $\sigma_{\delta}^{2}$, where $\sigma_{\delta}^{2}$ is assumed to have an inverse $\chi^{2}(1)$ prior. This reparmeterisation has no effect, of course, on the associated density: we retain the expression in terms of $\mu$ to allow readers to compare with our analogous expression for the distributional null.
The JZS Bayesian $t$-test is implemented in the R BayesFactor package via the function ‘ttest.tstat’ (Morey and Rouder,, 2021). To compare this approach against our distributional null proposal, we used this package to calculate $BF_{10}$ for every experiment in the Many-Labs dataset, and compared these $BF_{10}$ values against the distributional significance $p_{sig}$ for that result on a $\log_{10}$ scale (Figure 2(Right)). This comparison gave results almost identical those seen in the earlier point-form $p$ comparison. There were many extremely high $BF_{10}$ values, with $33\%$ of experiments having $BF_{10}>10^{5}$ and $17\%$ of experiments having $BF_{10}>10^{15}$. Distributional $p_{sig}$ values were again orders of magnitude more conservative. Many results gave strong or very strong evidence against the null in the Bayesian test while not reaching the minimum level of significance in the distributional test: $20\%$ of results with $BF_{10}>30$ (very strong evidence against the null) had $p_{sig}>0.05$. Even results with apparently very strong evidence against the null (in the Bayesian $t$-test) turn out to be consistent with the null hypothesis when between-experiment variation is taken into account.
The left and right scatterplots in Figure 2 are extremely similar, suggesting that the logs of the point-form $p$ and the Bayes Factor $BF_{10}$ are related; or more precisely (since Figure 2(Left) has a negative base-10 log scale on the $y$ axis) that $\log_{10}(BF_{10})$ and $\log_{10}(1/p)$ are related. A comparison of values of $\log_{10}(BF_{10})$ versus $\log_{10}(1/p)$ across experiments and tasks in the Many Labs dataset and confirms that these measures have an almost perfect linear relationship, with a correlation of $r=0.9999$ and with the line of best fit being $\log_{10}(BF_{10})=0.99\log_{10}(1/p)-1.59$. This means that the point-form and the Bayesian $t$-test convey essentially the same information, at least in this dataset. This close relationship between the point-form $p$ value and the Bayesian $t$-test measure is not surprising: it has long been observed empirically that Bayes Factor measures involving the point-form null and the point-form $p$-value measure are strongly correlated (e.g. Wetzels et al.,, 2011) and that decisions based on these two measures are often equivalent (to quote Jeffreys: “As a matter of fact I have applied my significance tests to numerous applications that have also been worked out by Fisher’s, and have not yet found a disagreement in the actual decisions reached”; cited in Ly et al.,, 2016). Given this relationship it seems clear that the various problems with point-form hypothesis testing (failures of replication for significant results, the fact that the point-form null hypotheses are always false, the effects of sample size on significance, and so on) will also apply to Bayesian approaches that involve the point-form null: we argue that these problems can be addressed, not by moving from a Frequentist to a Bayesian position, but instead by moving from point-form to distributional nulls.
6 Conclusions
The above results give evidence in favour of the distributional approach to significance testing, supporting its assumptions about Normally-distributed between-experiment variance of means, showing that many results that appear statistically significant under the point-form approach are identified as consistent with with random (between-experiment) variation in the distributional approach, and showing that the predicted replication probabilities generated for target experiments in the distributional approach are closely related to the actual observed probability of successful replication in those experiments, at least in the Many Labs 1 dataset.
We believe this distributional approach has a number of advantages over standard point-form significance testing. First, this distributional approach more accurately reflects the various random factors that do, in fact, influence all experimental outcomes and so gives a more accurate estimate of the degree to which a given outcome is unexpected or surprising under a null model of random variation but no real effect. Second, this approach addresses well-known problems with sample size that arise in the standard approach, where the probability of a significant result rises with sample size irrespective of any real effect. This issue is particularly important in research involving hypothesis testing on large data sets (increasingly common in the ‘big data’ era), and so the distributional approach should be useful in this form of research. Third, because the distributional model includes between-experiment variation in statistical analysis, it gives a natural account for the rate at which significant results will replicate in future experiments (an account which the point-form model cannot provide, because that model does not allow for cross-experiment variation in the experimental mean). Finally, this approach also gives a useful perspective on commonly used measures of the statistical power of an experiment to detect effects, showing that in the presence of between-experiment variation some effects simply cannot be detected reliably, no matter how large the sample.
We expect a number of objections to our argument for a distributional approach to hypothesis testing. The first objection concerns the use of a distributional representation of the null hypothesis. ‘These distributional nulls are just Bayesian priors in another form’ we imagine the objection goes, ‘and Bayesian priors are subjective measures of belief, not objective probability estimates. Subjective beliefs cannot enter into objective frequentist hypothesis testing’.
It is true that our distributional nulls have a mathematical form that is identical to a Bayesian prior. It is not true, however, that these distributional nulls necessarily represent subjective measures of belief: in a Frequentist interpretation, they can be taken to estimate between-experiment variance. Indeed, these distributional nulls play the same role that point-form nulls play in standard significance testing: for a given (distributional or point-form) null hypothesis we say ‘result $t$ would have a low probability of occurrence if null hypothesis were true, and so result $t$ is unlikely to be simply a consequence of random processes’. Such assertions do not require or reflect any subjective belief in the null hypothesis. To put this response another way: since the logic of significance testing is independent of the form of null hypothesis being used, our use of distributional rather than point nulls does not change the objective nature of such hypothesis testing.
A second objection concerns statistical testing against distributional null hypotheses characterised by the parameter $b$ (representing the between-experiment to within-experiment variance ratio). ‘Researchers can adjust the distribution of this parameter $b$ until they find a null hypothesis against which their observed results are statistically significant’ we imagine the objection goes. ‘But this is simply a form of data-dredging or p-hacking: an attempt to find patterns in data that can be presented as statistically significant when in fact there is no real underlying effect’.
Our response here is to point out that this value $b$ is not picked arbitrarily: instead, $b$ is estimated from experimental data on replications (and we give unbiased estimators for that situation). More generally, even if researchers were arbitrarily picking values for this parameter $b$, the use of distributional testing systematically reduces the occurrence of statistically significant results, relative to the point-form null. This is because the rejection region for a null hypothesis falls monotonically with the value of this ratio $b$: and so the maximum rejection region (and so the greatest chance of a statistically significant result) arises with the point-form null. A better characterisation of this process of testing against distributional nulls is one where we attempt to reduce the chance of finding statistically significant results: where we are conservative in results as statistically significant, taking into account both between-experiment and within-experiment variance.
A third possible objection concerns the availability of alternative hypothesis-testing methods, such as the Bayesian $t$-test (commonly put forward as a replacement for the significance testing approach). ‘While significance testing was appropriate in the last century’, we imagine the objection goes, ‘today we have better statistical methods; we don’t need to test for significance’. Our response here is to note that measures of significance in the point-form $t$-test and measures of evidence in the JZS Bayesian $t$-test appear to be almost exactly equivalent in order of magnitude, at least in the Many Labs $1$ data, and so replacing significance testing against a point-form null with this form of Bayesian $t$-test may not, therefore, represent a meaningful advance.
While useful, our distributional approach is subject to various caveats. First, our approach assumes normally distributed variation or error both within and between experiments. This assumption is reasonable for simple experiments involving one or two variables ($t$-tests, tests of correlation or contingency) but does not hold in more complex experimental designs involving multiple variables, such as ANOVAs, multiple regression analyses, and mixed models. This is because the internal structure in such designs can cause random between-experiment differences to have systematic non-normal effects on between-experiment variability. Applying our distributional approach to between-experiment variation in this more complex setting would be challenging, requiring a statistical model of between-experiment variance in each component of the experimental design and how that variance affects the overall between-experiment distribution of results.
Second, while our distributional approach allows us to estimate the probability of replication of given experimental results, it is not a ‘magic bullet’ that resolves all problems of replication. In particular, our approach does not address the important problem of publication bias. The distributional approach works best with some estimate for the between-experiment variability of means for a given experimental task, $S^{2}_{0}$. This can only be estimated from available (that is, published) data on experimental means. If this set of available means is restricted in some way as a consequence of selective publication (if, for example, published means tend to be further from 00 than unpublished means), this will lead to a systematic underestimation of between-experiment variation. This problem of publication bias does not apply to the Many Labs $1$ dataset results: while the experimental tasks examined in that dataset were selected because they investigate well-known classic and contemporary effects in psychology (and so the original papers describing those effects were subject to some degree of publication bias), the results from all $36$ replications of those effects were ‘published’ irrespective of their experimental means. This problem clearly does apply, however, if between-experiment variability is estimated from a set of individually published experimental results. This problem of publication bias is best addressed by continuing replication studies in various fields. Such studies are additionally important because they will give useful estimates of the degree of between-experiment variation in experiments of different types, allowing researchers to assess the approximate probability of replication of new results in similar experiments.